OpenStudy (anonymous):

You own 6 hats and are taking 4 on vacation. In how many ways can you choose 4 hats from the 6?

OpenStudy (anonymous):

nPr

OpenStudy (shubhamsrg):

I'd say nCr

OpenStudy (anonymous):

can you explain how to do this concept i'm still confused by it!

OpenStudy (mathstudent55):

You need to choose 4 out of 6, in a case in which order does not matter.

OpenStudy (anonymous):

the available answers are a) 360 b) 720 c) 1 D) 15

OpenStudy (mathstudent55):

For this, you need a combination, not a permutation. nCr = n!/((n-r)!r!) In your case, this is 6C4 = 6!/(6 - 4)!4!) = = (6 * 5 * 4!)/(2! * 4!) = = 6 * 5 / 2 = 15

OpenStudy (anonymous):

ok! thank you so much! thats makes more sense!!

OpenStudy (mathstudent55):

wlcm